OpenStudy (anonymous):

The Everton College store paid $1805 for an order of 47 calculators . The store paid $11 for each scientific calculator. The others, all graphing calculators, cost the store $57 each. How man scientific calculators were ordered? How many graphing calculators were ordered?

OpenStudy (anonymous):

let the no of scientifice calc = x then no of graph calc = 47 - x cost of x scient calc = 11x cost of 47- x graph calc = 57(47 - x) = 2679 - 57x total cost is 1805 => 11x + 2679 - 57x = 1805 => -46x = 1805-2679 => x = -874/-46 => x = 19 so no of scient calc = x = 19 no. of graph calc = 47 - x = 47 - 19 = 28

OpenStudy (anonymous):

Thank you... I always have ahard time with this.. I attempt to do simple math to figure out and it never works.

OpenStudy (anonymous):

Most welcome to my help...... if u r satisfied, pls click GOOD ANSWER button for me and also pls become my fan (:))

OpenStudy (anonymous):

I am your fan! =)

OpenStudy (anonymous):

hv u clicked the fan button too !!!

OpenStudy (anonymous):

ok i saw u hv....thx!!!

OpenStudy (anonymous):

I think so. I will dble check.

OpenStudy (anonymous):

I did.

OpenStudy (anonymous):

thx !!!!!